# Complexity of equal 0-surgeries

### v2

@@ 8. Extending diffeomorphisms over traces and symmetry-exceptional slopes
$K14n3611$ $\mathbb{Z}_{2}$ $-o9\_27542$ $\mathbb{Z}_{2}$ $JSJ\big{(}M\tilde{\times}S^{1}\cup m015\big{)}$ at least $\mathbb{Z}_{2}$
$K14n5084$ $\mathbb{Z}_{2}$ $-o9\_37547$ $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K15n19499$ $\mathbb{Z}_{2}$ $K15n153789$ 00 $JSJ\big{(}SFS\,[D^{2}:(2,1)(2,-1)]\cup m032\big{)}$ at least $\mathbb{Z}_{2}$
$K15n19499$ $\mathbb{Z}_{2}$ $o9\_31828$ $\mathbb{Z}_{2}$ $JSJ\big{(}SFS\,[D^{2}:(2,1)(2,-1)]\cup m032\big{)}$ at least $\mathbb{Z}_{2}$
$K15n94464$ 00 $-o9\_40081$ $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K15n101402$ 00 $-o9\_34908$ $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K15n103488$ $\mathbb{Z}_{2}$ $o9\_42735$ $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K15n153789$ 00 $o9\_31828$ $\mathbb{Z}_{2}$ $JSJ\big{(}SFS\,[D^{2}:(2,1)(2,-1)]\cup m032\big{)}$ at least $\mathbb{Z}_{2}$
@@
$v2869$ $\mathbb{Z}_{2}$ $t12388$ 00 hyperbolic $\mathbb{Z}_{2}$
$t07281$ $\mathbb{Z}_{2}$ $o9\_34949$ $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$t07281$ $\mathbb{Z}_{2}$ $o9\_39806$ $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$-t10974$ $\mathbb{Z}_{2}$ $o9\_39967$ $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}$
$t11462$ $\mathbb{Z}_{2}$ $-t11462$ $\mathbb{Z}_{2}$ hyperbolic $D_{4}$
$t11532$ 00 $-o9\_43446$ 00 hyperbolic 00
$t11900$ 00 $-o9\_40803$ 00 hyperbolic 00
$o9\_22951$ $\mathbb{Z}_{2}$ $-o9\_22951$ $\mathbb{Z}_{2}$ $JSJ\big{(}SFS\,[D^{2}:(3,1)(3,-1)]\cup m004\big{)}$ at least $\mathbb{Z}_{2}$
$o9\_39806$ $\mathbb{Z}_{2}$ $o9\_34949$ $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$o9\_41058$ 00 $-o9\_41058$ 00 hyperbolic $\mathbb{Z}_{2}$
@@ 9. Algorithmic search for 0-surgery diffeomorphisms
In this section, we explain and implement various methods to algorithmically search for a Kirby calculus presentation of a 00-surgery diffeomorphism of a given pair of friends. Using these methods we find explicit 00-surgery diffeomorphisms for many friends from the census of Table 1 and are in some cases able to answer if the friends are $4$-dimensional friends or not.
We also note here that there exists an algorithm that takes as input diagrams of a pair of friends $K$ and $K^{\prime}$ and outputs a 00-surgery diffeomorphism presented as a sequence of Kirby moves. Indeed, we can just enumerate all possible Kirby moves in the Kirby diagram of $K(0)$ given by a 00-framed diagram of $K$ and will reach in finite time to the Kirby diagram of $K^{\prime}(0)$ given by a 00-framed diagram of $K^{\prime}$. Of course, this algorithm is not practical but guaranteed to theoretically always work. Below we explain certain strategies that will not work in general but often return a verified result in reasonable time.
@@ 9.1. Piccirillo friends
Let $K$ be a knot with unknotting number $1$. Piccirillo [Pic20] describes a simple method for constructing a knot $K^{*}$, the Piccirillo friend, whose trace is diffeomorphic to the trace of $K$. Given a diagram $D$ of $K$ that realizes the unknotting number of $K$ this process is algorithmic. We have implemented that algorithm [AKW24] and have applied that code to prove the following.
@@
Of the knots in Table 1 we have found that at least $15$ knots have unknotting number one. For that, we load (potential) minimal crossing diagrams of these knots from [Bur20, BK24] and check which of these diagrams have unknotting crossings. We found that exactly $15$ of these diagrams admit an unknotting crossing. The other diagrams are proven to have no unknotting crossing and we expect the knots to have unknotting number larger than one.
Next, we implement the algorithm from [Pic20] and construct the Piccirillo friend $K^{*}$. Then we use SnapPy to check if the Piccirillo friend is isotopic to the friend in Table 1. This identifies that the above-claimed $9$ friends are $4$-dimensional friends. Note that this also yields a simple (human checkable) certificate that these knots share the same 00-surgeries.
In the other cases, the Piccirillo dual was a different knot than the one appearing in our census. ∎
9.2. Special RGB links
Manolescu–Piccirillo [MP23] describe a particularly simple family of RBG links, called special RBG links, with the property that $B$ and $G$ both are meridians of $R$. We transform a diagram of a special RBG link $L$ into a Kirby diagram by putting an integer framing $r$ on $R$ and framing 00 on both $G$ and $B$. Then we can perform a slam dunk to cancel $R$ and $B$. The image of $G$ under this cancellation is a knot $K_{G}$ in $S^{3}$. By symmetry, the same holds true with the roles of $G$ and $B$ reversed. By choosing the correct framing $r$ on $R$ we get two knots $K_{B}$ and $K_{G}$ that share the same 00-surgery.
Now the family of RBG links $L_{1}{(a,b,c,d,e,f)}$ from [MP23] shown in Figure 3(i) depends on six parameters $(a,b,c,d,e,f)$ and is chosen in such a way that the resulting knots $K_{B}$ and $K_{G}$ have relatively small crossing number. In Figure 3(ii) and (iii) we display two more similar RBG links $L_{2}{(a,b,c,d,e,f)}$ and $L_{3}{(a,b,c,d,e,f)}$. It is shown in [MP23] that the parity of $r=a+b$ yields the parity of the associated 00-surgery diffeomorphism (given by the slam dunks) and that if $r=0$ this 00-surgery diffeomorphism extends to a trace diffeomorphism. Below we use this family to find explicit Kirby moves for several pairs of friends from our census and study if they are $4$-dimensional friends or not.
@@ 9.2. Special RGB links
00
Figure 3. Three special RBG links. $L_{1}$ is the RBG link from Figure 13 in [MP23]. $L_{2}$ and $L_{3}$ are obtained from $L_{1}$ by crossing changes.
@@
The $16$ pairs of friends shown in Table 8 arise as knots $K_{B}$ and $K_{G}$ from the RGB link $L_{i}{(a,b,c,d,e,f)}$ for the printed values of ${(a,b,c,d,e,f)}$. The column ’traces’ describes how the traces of these friends are related.
Table 8. Friends obtained from special RGB links
@@
We import a surgery presentation of $L_{i}{(a,b,c,d,e,f)}$ into SnapPy which we can Dehn fill to obtain ideal triangulations of the complements of the associated knots $K_{B}$ and $K_{G}$ for small values of the parameters $(a,b,c,d,e,f)$. Then we check if SnapPy recognizes these triangulations as complements of knots in the census or in the low-crossing knots. Concretely, we check each of the parameters $(a,b,c,d,e,f)$ in the range $\{-10,\ldots,10\}$. Most pairs of friends that arise from $L_{i}{(a,b,c,d,e,f)}$ arise in several different ways. From the RBG link, we get an explicit sequence of Kirby moves describing a 00-surgery diffeomorphism. The parity of this diffeomorphism is the parity of $a+b$. In Table 8 we always list the simplest choice of parameters for each possible parity. If we found an even 00-surgery diffeomorphism we always tried to list one that has $a+b=0$, since for these we know that the 00-surgery diffeomorphism extends to a trace diffeomorphism. ∎
@@
For the friends where we only found odd 00-surgery diffeomorphisms, we can check from Table 6 that the 00-surgery is either exceptional or symmetry-exceptional and thus Proposition 8.3 cannot be applied. A possible approach to study these cases would be to compute the symmetry group $G$ of the 00-surgery. If $G$ is a finite group (for example if the 00-surgery is hyperbolic) we consider all possible compositions of the 00-surgery diffeomorphisms given by the special RBG links and see them as elements in $G$. Next, we follow the action of the meridians through these Kirby moves. From the action of the meridian under these compositions, we can distinguish the 00-surgery diffeomorphisms and find a composition of 00-surgery diffeomorphisms for every element in $G$. In particular, we can display also the exceptional symmetries like this. From the action on the meridians, we can also determine the parities of these 00-surgery diffeomorphisms.
@@ 9.3. Flat annulus presentations
A classical method to construct infinitely many mutually distinct knots that share the same 00-surgery is via annulus twisting [Oso06, AJOT13, AJLO15, AT21]. For examples, we refer to the two links $L_{1}$ and $L_{2}$ shown in Figure 4. Performing $(1/n)$- and $(-1/n)$-surgeries on the blue and red components yields a sequence of knots $K_{n}^{m}$. By the main result of [Oso06] it follows that for every $m\in\mathbb{Z}$ the knots $K_{n}^{m}$ are all friends. Moreover, such an annulus presentation yields an explicit sequence of Kirby moves describing 00-surgery diffeomorphisms $\varphi_{n}^{m}\colon K_{0}^{m}(0)\rightarrow K_{n}^{m}(0)$, see for example [Oso06]. If the core curve of the annulus is unknotted and if the boundary curves of the annulus have vanishing linking number (as the red and blue curves in Figure 4) it is straightforward to check that the parity of $\varphi_{n}^{m}$ is given by the parity of $n$ and that for any even $n$, the 00-surgery diffeomorphism $\varphi_{n}^{m}\colon K_{0}^{m}(0)\rightarrow K_{n}^{m}(0)$ extends to a trace diffeomorphism, see for example [MP23].
@@
The pairs of friends shown in Table 9 arise as knots $K_{n}^{m}$ from the flat annulus presentations shown in Figure 4. The column ’traces’ describes how the traces of these friends are related.
@@
We import the surgery descriptions $L_{1}$ and $L_{2}$ into SnapPy and use these to create ideal triangulations of the complements of $K_{n}^{m}$ for small values of $m$ and $n$. Whenever SnapPy recognizes a pair of these manifolds (with the same $m$ but different values of $n$) as manifolds from our census we print it in Table 9. By the above discussion, the parity of the 00-surgery diffeomorphism given by the Kirby moves induced from the annulus presentation is given by the parity of the difference of the values of $n$. If that difference is even, the traces are diffeomorphic and if it is odd, it follows that this particular 00-surgery diffeomorphism does not extend to a trace homeomorphism. In Examples 9.5 and 9.6 below, we use Proposition 8.3 to deduce that for some of these friends, no 00-surgery diffeomorphism extends to a trace homeomorphism. ∎
@@
From Table 6 we read-off that the three friends $(K10n10,t11532)$, $(K12n318,o9\_39433)$, and $(t11532,-o9\_43446)$ are all asymmetric and have asymmetric 00-surgery. Thus Proposition 8.3 implies that all 00-surgery diffeomorphisms have the same parity (since the 00-surgeries are asymmetric, all 00-surgery diffeomorphisms are even isotopic). From Table 9 we read off that these three pairs of friends each admit a 00-surgery diffeomorphism with odd parity and thus their traces cannot be homeomorphic.
We consider the pair of friends $K12n309$ and $-K14n14254$. Both knots are strongly invertible and have symmetry group $\mathbb{Z}_{2}$ generated by the strong inversion. In Table 6 we computed that the symmetry group of their 00-surgery is $\mathbb{Z}_{2}$, generated by the extension of the strong inversion. In particular, 00 is not symmetry-exceptional for both knots. By Proposition 8.3 all 00-surgery diffeomorphisms have the same pairity. Since the 00-surgery diffeomorphism from Table 9 is odd, Boyer’s theorem implies that no 00-surgery diffeomorphism of this pair extends to a homeomorphism of the traces.
@@
We consider the pair of friends $K13n469$ and $-K13n469$. The diffeomorphism of the 00-surgery obtained from the annulus twist in Table 9 does not extend to a diffeomorphism of the traces (since it has odd parity). However, the diffeomorphism of the 00-surgery from Remark 4.2 extends to a diffeomorphism of the traces. Thus Proposition 8.3 implies that 00 is an exceptional or symmetry-exceptional slope of $K13n469$. In fact, we can check that the 00-surgery of $K13n469$ is not hyperbolic, see Table 6.

### v3

@@ 8. Extending diffeomorphisms over traces and symmetry-exceptional slopes
$K14n3611$ $\mathbb{Z}_{2}$ o9_27542-o9\_27542 $\mathbb{Z}_{2}$ $JSJ\big(M\tilde{\times}S^{1}\cup m015\big)$ at least $\mathbb{Z}_{2}$
$K14n5084$ $\mathbb{Z}_{2}$ o9_37547-o9\_37547 $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K15n19499$ $\mathbb{Z}_{2}$ $K15n153789$ 0 $JSJ\big(SFS\,[D^{2}:(2,1)(2,-1)]\cup m032\big)$ at least $\mathbb{Z}_{2}$
$K15n19499$ $\mathbb{Z}_{2}$ o9_31828o9\_31828 $\mathbb{Z}_{2}$ $JSJ\big(SFS\,[D^{2}:(2,1)(2,-1)]\cup m032\big)$ at least $\mathbb{Z}_{2}$
$K15n94464$ 0 o9_40081-o9\_40081 $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K15n101402$ 0 o9_34908-o9\_34908 $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K15n103488$ $\mathbb{Z}_{2}$ o9_42735o9\_42735 $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K15n153789$ 0 o9_31828o9\_31828 $\mathbb{Z}_{2}$ $JSJ\big(SFS\,[D^{2}:(2,1)(2,-1)]\cup m032\big)$ at least $\mathbb{Z}_{2}$
@@
$v2869$ $\mathbb{Z}_{2}$ $t12388$ 0 hyperbolic $\mathbb{Z}_{2}$
$t07281$ $\mathbb{Z}_{2}$ o9_34949o9\_34949 $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$t07281$ $\mathbb{Z}_{2}$ o9_39806o9\_39806 $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
t10974-t10974 $\mathbb{Z}_{2}$ o9_39967o9\_39967 $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}$
$t11462$ $\mathbb{Z}_{2}$ t11462-t11462 $\mathbb{Z}_{2}$ hyperbolic $D_{4}$
$t11532$ 0 o9_43446-o9\_43446 0 hyperbolic 0
$t11900$ 0 o9_40803-o9\_40803 0 hyperbolic 0
o9_22951o9\_22951 $\mathbb{Z}_{2}$ o9_22951-o9\_22951 $\mathbb{Z}_{2}$ $JSJ\big(SFS\,[D^{2}:(3,1)(3,-1)]\cup m004\big)$ at least $\mathbb{Z}_{2}$
o9_39806o9\_39806 $\mathbb{Z}_{2}$ o9_34949o9\_34949 $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
o9_41058o9\_41058 0 o9_41058-o9\_41058 0 hyperbolic $\mathbb{Z}_{2}$
@@ 9. Algorithmic search for 0-surgery diffeomorphisms
In this section, we explain and implement various methods to algorithmically search for a Kirby calculus presentation of a 0-surgery diffeomorphism of a given pair of friends. Using these methods we find explicit 0-surgery diffeomorphisms for many friends from the census of Table 1 and are in some cases able to answer if the friends are $4$-dimensional friends or not.
We also note here that there exists an algorithm that takes as input diagrams of a pair of friends $K$ and $K^{\prime}$ and outputs a 0-surgery diffeomorphism presented as a sequence of Kirby moves. Indeed, we can just enumerate all possible Kirby moves in the Kirby diagram of $K(0)$ given by a 0-framed diagram of $K$. In finite time, we will reach the Kirby diagram of $K^{\prime}(0)$ given by a 0-framed diagram of $K^{\prime}$. Of course, this algorithm is not practical but guaranteed to theoretically always work. Below we explain certain strategies that will not work in general but often return a verified result in reasonable time.
@@ 9.1. Piccirillo friends
Let $K$ be a knot with unknotting number $1$. Piccirillo [47] describes a simple method for constructing a knot $K^{*}$, the Piccirillo friend, whose trace is diffeomorphic to the trace of $K$. Given a diagram $D$ of $K$ that realizes the unknotting number of $K$ this process is algorithmic. We have implemented that algorithm [3] and have applied that code to prove the following.
@@
Of the knots in Table 1 we have found that at least $15$ knots have unknotting number one. For that, we load (potential) minimal crossing diagrams of these knots from [17, 12] and check which of these diagrams have unknotting crossings. We found that exactly $15$ of these diagrams admit an unknotting crossing. The other diagrams are proven to have no unknotting crossing and we expect the knots to have unknotting number larger than one.
Next, we implement the algorithm from [47] and construct the Piccirillo friend $K^{*}$. Then we use SnapPy to check if the Piccirillo friend is isotopic to the friend in Table 1. This identifies that the above-claimed $9$ friends are $4$-dimensional friends. Note that this also yields a simple (human checkable) certificate that these knots share the same 0-surgeries.
In the other cases, the Piccirillo friend was a different knot than the one appearing in our census. ∎
9.2. Special RBG links
Manolescu–Piccirillo [39] describe a particularly simple family of RBG links, called special RBG links, with the property that $B$ and $G$ both are meridians of $R$. We transform a diagram of a special RBG link $L$ into a Kirby diagram by putting an integer framing $r$ on $R$ and framing 0 on both $G$ and $B$. Then we can perform a slam dunk to cancel $R$ and $B$. The image of $G$ under this cancellation is a knot $K_{G}$ in $S^{3}$. By symmetry, the same holds true with the roles of $G$ and $B$ reversed. By choosing the correct framing $r$ on $R$ we get two knots $K_{B}$ and $K_{G}$ that share the same 0-surgery.
Now the family of RBG links $L_{1}{(a,b,c,d,e,f)}$ from [39] shown in Figure 3(i) depends on six parameters $(a,b,c,d,e,f)$ and is chosen in such a way that the resulting knots $K_{B}$ and $K_{G}$ have relatively small crossing number. In Figure 3(ii) and (iii) we display two more similar RBG links $L_{2}{(a,b,c,d,e,f)}$ and $L_{3}{(a,b,c,d,e,f)}$. It is shown in [39] that the parity of $r=a+b$ yields the parity of the associated 0-surgery diffeomorphism (given by the slam dunks) and that if $r=0$ this 0-surgery diffeomorphism extends to a trace diffeomorphism. Below we use this family to find explicit Kirby moves for several pairs of friends from our census and study if they are $4$-dimensional friends or not.
@@ 9.2. Special RBG links
0
Figure 3. Three special RBG links. $L_{1}$ is the RBG link from Figure 13 in [39]. $L_{2}$ and $L_{3}$ are obtained from $L_{1}$ by crossing changes.
@@
The $16$ pairs of friends shown in Table 8 arise as knots $K_{B}$ and $K_{G}$ from the RBG link $L_{i}{(a,b,c,d,e,f)}$ for the printed values of (a,b,c,d,e,f){(a,b,c,d,e,f)}. The column ’traces’ describes how the traces of these friends are related.
Table 8. Friends obtained from special RBG links
@@
We import a surgery presentation of $L_{i}{(a,b,c,d,e,f)}$ into SnapPy which we can Dehn fill to obtain ideal triangulations of the complements of the associated knots $K_{B}$ and $K_{G}$ for small values of the parameters $(a,b,c,d,e,f)$. Then we check if SnapPy recognizes these triangulations as complements of knots in the census or in the low-crossing knots. Concretely, we check each of the parameters $(a,b,c,d,e,f)$ in the range $\{-10,\ldots,10\}$. Most pairs of friends that arise from $L_{i}{(a,b,c,d,e,f)}$ arise in several different ways. From the RBG link, we get an explicit sequence of Kirby moves describing a 0-surgery diffeomorphism. The parity of this diffeomorphism is the parity of $a+b$. In Table 8 we always list the simplest choice of parameters for each possible parity. If we found an even 0-surgery diffeomorphism we always tried to list one that has $a+b=0$, since for these we know that the 0-surgery diffeomorphism extends to a trace diffeomorphism. ∎
@@
For the friends where we only found odd 0-surgery diffeomorphisms, we can check from Table 6 that the 0-surgery is either exceptional or symmetry-exceptional and thus Proposition 8.3 cannot be applied. A possible approach to study these cases would be to compute the symmetry group $G$ of the 0-surgery. If $G$ is a finite group (for example if the 0-surgery is hyperbolic) we consider all possible compositions of the 0-surgery diffeomorphisms given by the special RBG links and see them as elements in $G$. Next, we follow the action of the meridians through these Kirby moves. From the action of the meridian under these compositions, we can distinguish the 0-surgery diffeomorphisms and find a composition of 0-surgery diffeomorphisms for every element in $G$. In particular, we can display also the exceptional symmetries like this. From the action on the meridians, we can also determine the parities of these 0-surgery diffeomorphisms.
@@ 9.3. Flat annulus presentations
A classical method to construct infinitely many mutually distinct knots that share the same 0-surgery is via annulus twisting [43, 2, 1, 4]. For examples, we refer to the two links $L_{1}$ and $L_{2}$ shown in Figure 4. Performing $(1/n)$- and (1/n)(-1/n)-surgeries on the blue and red components yields a sequence of knots $K_{n}^{m}$. By the main result of [43] it follows that for every $m\in\mathbb{Z}$ the knots $K_{n}^{m}$ are all friends. Moreover, such an annulus presentation yields an explicit sequence of Kirby moves describing 0-surgery diffeomorphisms $\varphi_{n}^{m}\colon K_{0}^{m}(0)\rightarrow K_{n}^{m}(0)$, see for example [43]. If the core curve of the annulus is unknotted and if the boundary curves of the annulus have vanishing linking number (as the red and blue curves in Figure 4) it is straightforward to check that the parity of $\varphi_{n}^{m}$ is given by the parity of $n$ and that for any even $n$, the 0-surgery diffeomorphism $\varphi_{n}^{m}\colon K_{0}^{m}(0)\rightarrow K_{n}^{m}(0)$ extends to a trace diffeomorphism, see for example [39].
@@
The pairs of friends shown in Table 9 arise as knots $K_{n}^{m}$ from the flat annulus presentations shown in Figure 4. The column ‘traces’ describes how the traces of these friends are related.
@@
We import the surgery descriptions $L_{1}$ and $L_{2}$ into SnapPy and use these to create ideal triangulations of the complements of $K_{n}^{m}$ for small values of $m$ and $n$. (Concretely, we have checked $n,m\in\{-10,10\}$.) Whenever SnapPy recognizes a pair of these manifolds (with the same $m$ but different values of $n$) as manifolds from our census, we print it in Table 9. By the above discussion, the parity of the 0-surgery diffeomorphism given by the Kirby moves induced from the annulus presentation is given by the parity of the difference of the values of $n$. If that difference is even, the traces are diffeomorphic and if it is odd, it follows that this particular 0-surgery diffeomorphism does not extend to a trace homeomorphism. In Examples 9.5 and 9.6 below, we use Proposition 8.3 to deduce that for some of these friends, no 0-surgery diffeomorphism extends to a trace homeomorphism. ∎
@@
From Table 6 we read off that the three pairs of friends
$(K10n10,t11532),\,(K12n318,o9\_39433),\,\textrm{ and }\,(t11532,-o9\_43446)$
are all asymmetric and have asymmetric 0-surgery. Thus Proposition 8.3 implies that all 0-surgery diffeomorphisms have the same parity (since the 0-surgeries are asymmetric, all 0-surgery diffeomorphisms are even isotopic). From Table 9 we read off that these three pairs of friends each admit a 0-surgery diffeomorphism with odd parity and thus their traces cannot be homeomorphic.
We consider the pair of friends $K12n309$ and K14n14254-K14n14254. Both knots are strongly invertible and have symmetry group $\mathbb{Z}_{2}$ generated by the strong inversion. In Table 6 we computed that the symmetry group of their 0-surgery is $\mathbb{Z}_{2}$, generated by the extension of the strong inversion. In particular, 0 is not symmetry-exceptional for both knots. By Proposition 8.3 all 0-surgery diffeomorphisms have the same pairity. Since the 0-surgery diffeomorphism from Table 9 is odd, Boyer’s theorem (Theorem 8.1) implies that no 0-surgery diffeomorphism of this pair extends to a homeomorphism of the traces.
@@
We consider the pair of friends $K13n469$ and K13n469-K13n469. The diffeomorphism of the 0-surgery obtained from the annulus twist in Table 9 does not extend to a diffeomorphism of the traces (since it has odd parity). However, the diffeomorphism of the 0-surgery from Remark 4.2 extends to a diffeomorphism of the traces. Thus Proposition 8.3 implies that 0 is an exceptional or symmetry-exceptional slope of $K13n469$. In fact, we can check that the 0-surgery of $K13n469$ is not hyperbolic, see Table 6.